FLAO1 on telescope. Detailed Log of activities

FLAO1 run #1

FLAO1 run #2

FLAO1 run #3

FLAO1 run #4

  • LabLog20101013 bin1 con vibrazioni
  • LabLog20101014 bin1 con tree house spenta, bin2, bin3
  • LabLog20101015 KL_v6 400modi Bin1, Bin3, Bin4, KL_v6 500modi bin1, KL_v7 500modi bin 1 3l/d
  • LabLog20101019 KL_v6 bin1 3l/d, KL_v7 bin1 2l/d, gain automatico, KL_v6 bin2
  • LabLog20101020 Test astigmatismo, Misura di rumore IM, Modifica generazione disturbo, Modifica ampiezze, REC KL_v6 bin1 3l/d bin1 2l/d, bin2, bin3, bin4
  • LabLog20101021 Vignett filo alimentatore FW fioriera, test software con iif, Riproducibilta' ricostruttore Kl_v6 bin1 con dubbi su SR vs (pos psf on irtc)
  • LabLog20101022 Robustezza rec 10modi, KL_v7 bin1 3l/d bin1 2l/d, bin2, bin3, bin4, Open loop transfer function modo 7
  • LabLog20101023 Rotazione bin2 a passi di 5gradi, KL_v7 bin1 1l/d
  • LabLog20101024
  • LabLog20101025
  • LabLog20101026 Dome closed - Blade Runner
  • LabLog20101027 On Sky! AO738, AO730, AO56, AO960, [AO74, AO97, AO195], AO80, AO838, AO940, AO182 [IRTC FUOCO SBAGLIATO]
  • LabLog20101028 AO741, AO793, AO743, AO735, AO1066, AO802, AO1067, AO128, HD15745, LKHA330, AO1072, AO1073, AO849

FLAO1 run #5

  • LabLog20101115 Daytime: functionality check, cloop and first measurment of injected defocus
  • LabLog20101116 Daytime: test del fuoco modulato e testa di moro a n modi differenti
  • LabLog20101117 Daytime: SWEEPGAIN e testa di moro
  • LabLog20101118 Daytime: Test testa di moro e check movimento pupille con derot a tutti i binning
  • LabLog20101119 Daytime: testa di moro risolta;rivisitazione della tabella
  • LabLog20101120 Daytime: rivisitazione tabella delle configurazioni. Mettiamo la ND per andare faint
  • LabLog20101121 Daytime: test flat, rivisitazione tablle bin2 e bin3, fuoco IRTC daytime con ND. Nightime cloudy: dome closed, retro ON
  • LabLog20101122 Night: AO738 R7, AO49 R7, OC_M103 offaxis test, OC_m36 offaxis test, AO171 R7, AO263 R10.5
  • LabLog20101123 Daytime: retro off. Nightime windy and cloudy: dome closed, retro OF
  • LabLog20101124 Night: aperti solo 1h, problema LC x BCU47 troppo fredda
  • LabLog20101125 Night: AO794, AO704 acq low elev, AO777, AO74, AO78, LKHA330, AO247, AO327
  • LabLog20101126 Night:
  • LabLog20101127 Night:

FLAO1 run #6

FLAO1 run #7 (may-June 2011)

FLAO1 run #8 (September 2011)

FLAO1 run #9 (October 2011)

  • FLAO1LabLog20111008 Daytime: Startup FLAO1, adsecdx. Nighttime: test AO without PISCES, some TV images
  • FLAO1LabLog20111009 Daytime: Retroreflector
  • FLAO1LabLog20111010 Daytime: Retroreflector - autogain, optical loop gain. Fix 2 actuators. Night time: test AO without PISCES, some autogain and optloopgain.
  • FLAO1LabLog20111011 Daytime: Retroreflector - autogain. Night time: preset and autogain bin 1 bin2, offsets
  • FLAO1LabLog20111012 Night time: preset autogain bin 3, standards, off-axis reference star fail

SDT Run PISCES

* FLAO1LabLog20111108science SKY Ks, H, J

AO 28-29 Nov IT Run

FLAO2 Run#1 Daytime December 2011

FLAO2 Run#2 Jan-Feb 2012

FLAO2 Run#3 Mar-Apr 2012

FLAO2 Run#4 Apr-May 2012

FLAO2 Run#5 May-June 2012

INAF PISCES JUNE

FLAO2 Run#6 Sept 2012

FLAO1 run #10 (REMOTE)

FLAO2 REMOTE test cameralens

FLAO2LabLog20121123 Remote daytime

FLAO#1 run 2013-01

Dal 16 al 19/01 lavoro in clean room per sostituzione BCU47 e altri interventi. Vedi report in DOCX di Simone

20/01 AGW#2 rimontato sul FBG DX

FLAO#1&2 run 2013-03 remote support

FLAO#1 run 2013-08 with TS4

FLAO#1 run PISCES 2013-10 with TS4-DX

SHARK Forerunner 2014-02 on LBTI DX

FLAO#1 action during ARGOS run

FLAO#1 2014 October Run. Alignment to LUCI window (IRTC)

FLAO#1+LUCI2 Commissioning run #1 January 2015

SHARK Forerunner 2015-02 on LBTI DX

FLAO#1+LUCI2 Commissioning run #2 March 2015

FLAO#1+LUCI2 Commissioning run #3 May 2015 (remote support)

LBTI remote support May 2015

FLAO#1 test after BCU repair May 2015

FLAO#1 run May 2015

SHARK Forerunner 2015-06 on LBTI DX

SHARK Simulazioni

FLAO#2 (SX) alignment Feb 2016

FLAO#2 Daytime with LUCI1 March 2016

FLAO2TelLog20160312 NCPA Rot 330 and 300

FLAO2TelLog20160313 NCPA rot 60 120 and 150 deg. Test SW

FLAO2TelLog20160314 NCPA measurement and SW test

FLAO2TelLog20160315

FLAO2TelLog20160316 Failure GOPT

FLAO2TelLog20160317 FIRST LIGHT on sky

FLAO2TelLog20160318

FLAO#2 Nighttime witht LUCI1 April 2016

FLAO2TelLog20160421 (22 in Italia, supporto remoto)

FLAO2TelLog20160422 (23 in Italia, supporto remoto) WorkAround WorkAround2

2016B FLAO checkout by D. Miller

FLAOTelLog20160910
Topic revision: r265 - 12 Sep 2016, EnricoPinna
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback